OpenStudy (luigi0210):

Integrate:

OpenStudy (luigi0210):

\[\LARGE \int \frac{ln(x)}{x^2}~dx\]

OpenStudy (akashdeepdeb):

Try integration by parts. \[\int\limits_{}^{} u.dv = uv - \int\limits_{}^{}v.du\] Take u = ln (x) du = dx/x dv = x^(-2) . dx v = -x^(-1) That'll give you: \[- \ln (x)/x + \int\limits_{}^{} x^{-2}.dx\]

OpenStudy (akashdeepdeb):

And then integrate the second term, you'll have the answer.

OpenStudy (luigi0210):

Alright, thank you :)